# nnMamba: 3D Biomedical Image Segmentation, Classification and Landmark Detection with State Space Model

### 05 Feb 2024 v1

@@ 4 Conclusion
In summary, our work has introduced nnMamba, a novel framework that effectively fuses the granular feature extraction capabilities of CNNs with the expansive long-range dependency modeling afforded by State Space Sequence Models (SSMs). By innovatively integrating SSMs within the convolutional residual blocks, nnMamba has been tailored to excel in the domain of 3D medical image analysis, tackling image segmentation, classification, and landmark detection with unprecedented efficacy.
Our extensive experimental analysis confirms that nnMamba sets a new benchmark, surpassing existing state-of-the-art methodologies in handling complex image analysis tasks. The architecture’s capacity to seamlessly blend attention to local details with an understanding of global context makes it an essential tool in the field. The implications of nnMamba’s performance are substantial, offering a robust and versatile framework that promises to elevate the standard of medical image analysis and support the advancement of medical diagnostic and therapeutic applications.
It should be noted that current version only provides primary results. We are still working on this project by evaluating more backbones and datasets.
@@
This work was supported in part by the National Natural Science Foundation of China (NO. 62102267), and in part by the Natural Science Foundation of Guangdong Province of China (2023A1515011464).

### 10 Mar 2024 v2

@@ 4 Conclusion
Our work presents nnMamba, a new framework combining the detailed feature extraction of CNNs with the broad dependency modeling of SSMs. Designed for 3D medical image analysis, nnMamba excels in segmentation, classification, and landmark detection. It outperforms current methodologies, offering a blend of local and global image context understanding. nnMamba’s robust performance has significant implications for improving medical diagnostics and therapies.
@@
This work was supported in part by the National Natural Science Foundation of China (NO. 62102267), and in part by the Guangdong Basic and Applied Basic Research Foundation (2023A1515011464).